OpenStudy (shepherdll):

Find Lim( (sin3x/x) -1) as x --> 0

zepdrix (zepdrix):

Hey, welcome to OpenStudy! :) This is our problem? \[\large\rm \lim_{x\to0}~\frac{\sin(3x)}{x}-1\]

OpenStudy (shepherdll):

Yea that is the problem! My answer was 2 but it was wrong so i dont know what to do.

zepdrix (zepdrix):

This is a special limit identity that you learn early on in Calculus,\[\large\rm \lim_{y\to0}\frac{\sin(\color{orangered}{y})}{\color{orangered}{y}}\quad=\quad 1\]The important thing for this identity is that the angle matches the denominator.

zepdrix (zepdrix):

So for your problem, you need to find a fancy way to get a 3 in the denominator. You want to create a 3x in the denominator so it matches the argument of the sine.

zepdrix (zepdrix):

Wait... you ended up with 2? That shouldn't be wrong D: Hmm

OpenStudy (shepherdll):

Yea thats what i was thinking. My professor probably made a mistake than.

zepdrix (zepdrix):

Ya, weird :)